OpenStudy (lukecrayonz):

4=sqrt(6x+1)-sqrt(17-4x) 4=sqrt(6x+1)-sqrt(17-4x) @Mathematics

OpenStudy (anonymous):

4=√[(6x+1)-(17-4x)] (this is just another way to write it) now take away the √ from the right by taking 4^2 16=(6x+1)-(17-4x) now distribute your negative sign 16=6x+1-17+4x put together like terms 16=10x-16 and solve 32=10x > x=32/10 simplify fraction x=16/5

OpenStudy (lukecrayonz):

Its not one big sqrt, its two different ones, subtracted

OpenStudy (anonymous):

two square roots on one side of the equation are equivalent to one square root. the only thing on that side that you aren't taking the square root of is the negative sign and that doesn't make a difference to the total value you obtain for x. the only time you need to worry about solving for two individual square roots on one side of an equation is if you were dividing them or multiplying them

OpenStudy (lukecrayonz):

Well i used an online calculator and the answer was 4..

OpenStudy (lukecrayonz):

http://openstudy.com/#/updates/4ea471fae4b09b57843151c2 There goes someone else answering my question but stopped at the end, so if that helps, yeah :P

OpenStudy (anonymous):

okay i looked at the last thing they posted and its going to be very difficult foiling. has you math class learned how to foil 25x^2-56x-128. i can do it but it won't turn out to be 4

OpenStudy (lukecrayonz):

yes

OpenStudy (anonymous):

srry but i don't see how you can foil 25x^2-56x-128.

OpenStudy (anonymous):

4^2 [(√6x+1)-(√17-4x)]^2 16 = [(√6x+1)-(√17-4x)] * [(√6x+1)-(√17-4x)] solve this and it will get you 4. srry i couldn't solve this out for you but it is super long. just make sure to look at all of your rules :) good luck